Solaimon:

Question Toyota Prius (2021) Honda Accord (2020) Vehicle Make/Model/Year 2021 Toyota Prius (Hybrid) 2020 Honda Accord (Gasoline) Sales Price ($) 28000 24000 EPA Fuel Economy Rating (MPG/MPGe) 56 30 Sales Tax Rate (%) 4.15 4.15 Total Cost (Price + Tax) 29162 24996 Down Payment (20%) 5600 4800 Loan Amount 23562 20196 Loan Term (Months) 60 60 Loan Interest Rate (%) 5 5 Monthly Loan Payment ($) 444.6440071 381.1234347 Total Loan Cost (including Down Payment) 32278.64043 27667.40608 Monthly Insurance Cost ($) 75 75 Monthly Maintenance Cost ($) 100 100 Monthly Registration Cost ($) 2.5625 2.5625 Monthly Property Tax Cost ($) 100.3333333 86 Monthly Fuel Costs ($) 62.5 116.6666667 Total Monthly Ownership Cost ($) 785.0398405 761.3526013 How To answer for this using excel formulas?

Solaimon:

1 attachment
Solaimon:

Question is what excel formula would I use for this?

Midnight97:

@solaimon wrote:
Question is what excel formula would I use for this?
Do you just need the formula?

Solaimon:

Yes an excel formula

toga:

To calculate the costs for the Toyota Prius and Honda Accord in Excel, follow these steps: 1. Sales Price: Enter prices in cells A2 (Prius) and B2 (Accord). 2. Total Cost (Price + Tax): Use `=A2 * (1 + C3/100)` for Prius and `=B2 * (1 + C3/100)` for Accord, where C3 is the tax rate. 3. Down Payment (20%): Calculate as `=A2 * 0.20` for Prius and `=B2 * 0.20` for Accord. 4. Loan Amount: Subtract down payments from total costs: `=D2 - D5` (Prius) and `=D3 - D6` (Accord). 5. Monthly Loan Payment: Use PMT function: `=PMT(E5/12/100, F5, -G5)` for both. 6. Total Loan Cost: Calculate as `=(H5 * F5) + G5`. 7. Monthly Fuel Costs: For Prius, use `=(Average_miles_driven / 56) * Fuel_price` and do similarly for Accord. 8. Monthly Property Tax: Manually input values as needed. 9. Total Monthly Ownership Cost: Sum all monthly expenses: `=H5 + I5 + J5 + K5 + L5 + M5` for Prius and `=H6 + I6 + J6 + K6 + L6 + M6` for Accord.